OpenStudy (anonymous):

Misha used a probability simulator to pull 3 colored marbles from a bag and flip a coin 50 times. The results are shown in the tables below: Color of Marble Number of Times Rolled Blue 20 Green 18 Yellow 12 Heads Tails 28 22 Using Misha's simulation, what is the probability of pulling a blue marble and the coin landing tails up? 410 over 2500 440 over 2500 32 over 50 42 over 50

OpenStudy (kropot72):

The probability of selecting a blue marble is 20/50. The probability of tails is 22/50. The events are independent, therefore the probability of pulling a blue marble and the coin landing tails up is \[\large P(blue \cap tails)=\frac{20}{50}\times\frac{22}{50}=you\ can\ calculate\]

OpenStudy (anonymous):

0.176 is??? @kropot72

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

Although your calculation is correct, the choices of answer do not include this simplified result. Therefore you should not simplify, which gives: \[\large P(blue \cap tails)=\frac{20}{50}\times\frac{22}{50}=\frac{440}{2500}\]